OpenStudy (anonymous):

Solve the inequality: -7x + 4 ≤ 32

OpenStudy (kinggeorge):

\[-7x+4 \leq 32 \rightarrow -7x \leq 28 \rightarrow 7x \geq -28 \rightarrow x \geq -4\] so the ultimate solution is \[x \geq -4\]

OpenStudy (anonymous):

subtract oth sides you are left with -7x\[\le\]28 now divide both side by -7 x now eqauls -4